# DVPO: Distributional Value Modeling-based Policy Optimization for LLM Post-Training

### v1

@@ C.1.1 Dataset Construction
C.1.2 Training Configuration
C.1.3 Dialogue Task Evaluation
To further ensure reliability, we conducted cross-validation through two key steps: 1) multiple sampling runs to verify the stability of rubric execution, and 2) spot-checked human evaluations that demonstrated high agreement with rubric-based scores. To illustrate the evaluation logic for dialogue performance, the core prompt section regarding the assessment of conversation logicality is provided as follows in Figure 6, which is a part of goal completion rate.
Core Prompt for Dialogue Logicality When scoring the intelligent call assistant’s responses, please follow these steps: 1.Read and understand the entire call content. 2.Score from the dimension of dialogue logic rationality. The definition of dialogue logic rationality is: whether all responses provided by the assistant during the entire dialogue are clear, coherent, and effectively convey relevant information. A reasonable response should ensure the consistency, coherence, and relevance of information. Use a 1-5 scoring scale with specific criteria as follows (Due to space limitations, here we provide the end-point prompt): 1 point: The assistant’s responses have obvious logical contradictions multiple times. The responses violate basic logical rules. The user cannot understand the assistant’s intention, and the entire dialogue contains no useful information. 5 points: All of the assistant’s responses are completely consistent with rigorous logic, clear and effective information transmission, and very smooth overall dialogue; the responses are not only relevant but also can guide the in-depth development of the dialogue, promote better understanding and communication, and improve the user experience. The assistant takes the initiative to ask for relevant information and guides the end of the call. 3.Based on the above scoring criteria and combined with the entire dialogue content, give a reasonable score, along with the scoring thinking process, deduction points, and modification suggestions to improve the score without changing the original meaning of the text.
C.2 Additional experimental details in the Dialogue Task
@@
Table 7 presents the comprehensive results across all five real-world dialogue domains. DVPO achieves the highest overall performance with an average score of 86.79%, outperforming all RL baselines. Specifically, DVPO surpasses PPO and Reinforce++ by 1.59% and 2.94% percentage points on average, and exceeds GRPO-based methods by more than 20 percentage points in most domains. This advantage is consistent across all metrics: task completion, ask completion, and goal completion.
C.3 Additional Ablation Experiment
@@
Base 2.89% 3.10% 2.89% 87.40% 41.67% 28.68% 75.83% 2.96% 58.40% 34.64%
@@
Effectiveness of Calibrating Risk-Aware Distributional Optimization
Table 8 reports the ablation study on the distributional loss components in DVPO. Beginning from the core quantile regression baseline with an average accuracy of 36.63%, each additional objective provides a clear performance gain, showing that the losses capture complementary aspects of distributional modeling. Tail Calibration brings the first substantial improvement, raising out-of-distribution accuracy from 61.37% to 64.42% and improving the overall score to 38.42%. Shift Penalization further enhances robustness, especially under out-of-distribution conditions, achieving an accuracy of 65.17% in this setting. The combination of Tail Shape and Curvature yields the most balanced progress across both in-domain and out-of-domain tasks, ultimately reaching the highest average accuracy of 39.63%. These results demonstrate that distributional consistency, tail-aware calibration, and shape-sensitive constraints work synergistically to strengthen DVPO’s robustness and generalization under noisy supervision.
@@ C.4 Additional Visualization
Figures 7, 8, and 9 illustrate a comparative analysis of advantage estimations among PPO, Robust Bellman PPO, and our proposed method (DVPO) for the same question outputs, irrespective of their semantic quality.
Notably, significant discrepancies in advantage estimation patterns are observed across the three methods. For PPO, despite correctly identifying the final output as erroneous, the advantage values assigned to individual tokens exhibit minimal variability. Similarly, Robust Bellman PPO fails to recognize the invalidity of the novel output and also demonstrates negligible differences in token-level advantage estimations. Additionally, it overestimates the advantage of the final output component. This observation partially reflects the limited generalization capability of these two methods when handling novel response generations.

### v2

@@ C.1.1 Dataset Construction
Scenario Wealth Rental Insurance Food Express Promotion Loan Housing Service Product General Avg
Dialogue Count 87 99 138 120 215 66 70 87 67 69 92 94.73
Avg Turns 5.40 4.22 5.22 3.37 3.76 4.55 4.33 4.68 5.46 4.44 3.46 4.44
Table 6: Dataset statistics across 11 real-world service scenarios. The validation set covers diverse interaction types with varying dialogue lengths, enabling reliable evaluation of robustness and generalization under realistic settings.
Category Human Annotators GPT Models Human AVG Model AVG
Score 4.595 4.689 4.490 4.330 4.510 4.501 4.526 4.505
Table 7: Comparison of goal completion rate scores between human annotators and GPT-4o. Human annotators achieve an average score of 4.526, while GPT-4o models reach 4.505, showing a negligible performance gap.
@@
Base 1.78 2.36 2.87 84.80 35.00 22.43 75.00 2.55 54.31 32.13
GRPO 1.60 1.99 3.94 82.20 33.33 21.32 77.50 2.51 53.59 31.70
PPO 1.91 2.54 3.94 83.80 33.33 22.43 79.17 2.80 54.68 32.45
Reinforce++ 1.48 2.36 3.52 84.60 38.33 23.16 75.83 2.45 55.48 32.75
Dr.GRPO 1.50 2.17 3.10 84.00 35.00 22.79 76.67 2.26 54.62 32.18
Robust Bellman 1.64 2.17 4.17 84.60 31.67 22.79 75.00 2.66 53.52 31.72
DVPO (Ours) 1.73 2.36 4.36 86.20 43.33 25.00 78.33 2.82 58.22 34.47
Domain In-Domain (Science & QA) Out-of-Domain (Math) ID OOD ALL
Base 3.12% 3.80% 2.89% 88.40% 50.00% 31.25% 80.83% 3.27% 62.62% 37.18%
@@
C.1.2 Training Configuration
C.1.3 Dialogue Task Evaluation
To further ensure reliability, we conducted cross-validation through two key steps: 1) multiple sampling runs to verify the stability of rubric execution, and 2) spot-checked human evaluations that demonstrated high agreement with rubric-based scores. Specifically, for the comparative experiment between human annotators and the model, we validated the protocol using 1,110 data samples covering 10 scenarios. The statistical details of the validation dataset are presented in Table 6. Four independent annotators scored these samples strictly in accordance with the rubric. They are professional data annotators in the company. Table 7 presents the experimental results for goal completion rate scoring (full score: 5) in dialogue tasks. As shown in Table 7, the mean score assigned by human annotators is extremely close to that of the model, with a difference of only 0.021 points. These findings confirm strong consistency between human and model scores, validating the effectiveness of our rubric. We acknowledge that fully establishing external validity necessitates additional independent annotators. We are in the process of releasing our rubric, evaluation prompts, and real-dialogue dataset to facilitate replication. To illustrate the evaluation logic for dialogue performance, the core prompt section regarding the assessment of conversation logicality is provided as follows in Figure6, which is a part of goal completion rate.
Core Prompt for Dialogue Logicality When scoring the intelligent call assistant’s responses, please follow these steps: 1.Read and understand the entire call content. 2.Score from the dimension of dialogue logic rationality. The definition of dialogue logic rationality is: whether all responses provided by the assistant during the entire dialogue are clear, coherent, and effectively convey relevant information. A reasonable response should ensure the consistency, coherence, and relevance of information. Use a 1-5 scoring scale with specific criteria as follows: 1 point: The response exhibits severe logical inconsistencies and violates fundamental reasoning principles. The intent of the assistant is unclear, and the response fails to convey meaningful or usable information. 2 points: The response contains multiple inconsistencies or conflicting statements. Although partially related to the query, it lacks sufficient contextual grounding and coherent reasoning, resulting in poor interpretability and weak informational value. 3 points: The response is generally coherent but contains minor logical gaps or discontinuities that affect fluency. While relevant to the query, the explanation lacks depth or clarity in parts, limiting the overall effectiveness of communication. 4 points: The response is logically consistent and well-structured, with clear and relevant content. Information is conveyed effectively and aligns well with the user’s intent, enabling smooth and coherent interaction. 5 points: The response demonstrates strong logical consistency, clear reasoning, and precise information delivery. It not only addresses the query accurately but also facilitates deeper understanding, effectively guiding the conversation and enhancing overall interaction quality. 3.Based on the above scoring criteria and combined with the entire dialogue content, give a reasonable score, along with the scoring thinking process, deduction points, and modification suggestions to improve the score without changing the original meaning of the text.
C.2 Additional experimental details in the Dialogue Task
Table 8 presents the comprehensive results across all five real-world dialogue domains. DVPO achieves the highest overall performance with an average score of 86.79%, outperforming all RL baselines. Specifically, DVPO surpasses PPO and Reinforce++ by 1.59% and 2.94% on average, and exceeds GRPO-based methods by more than 20 percentage points in most domains. This advantage is consistent across all metrics: task completion, ask completion, and goal completion.
C.3 Additional Ablation Experiment
DVPO exhibits strong scalability across model scales
Effectiveness of Distributional Constraints on Robustness and Generalization
Table 10 presents an ablation study of different distributional loss components in DVPO.Starting from the core quantile regression baseline, we observe that introducing additional distributional constraints consistently improves performance across most settings. Among all components, tail-related objectives contribute the most significant gains. While distribution consistency and tail calibration improve robustness to some extent, their effects are more pronounced on out-of-domain tasks, especially on challenging math benchmarks such as AIME24 and AMC23. Notably, incorporating tail shape and curvature modeling leads to the most consistent improvements across both in-domain and out-of-domain evaluations, achieving the highest overall accuracy of 39.63%. These results indicate that explicitly modeling the shape of the value distribution plays a critical role in improving generalization. Compared to variance-reduction or calibration-based objectives alone, shape-aware constraints provide stronger structural regularization, enabling DVPO to better capture distributional properties under noisy or shifted reward settings.
@@ C.4 Additional Visualization
Figures 7, 8, and 9 illustrate a comparative analysis of advantage estimations among PPO, Robust Bellman PPO, and our proposed method (DVPO) for the same question outputs, irrespective of their semantic quality. Notably, significant discrepancies in advantage estimation patterns are observed across the three methods. For PPO, despite correctly identifying the final output as erroneous, the advantage values assigned to individual tokens exhibit minimal variability. Similarly, Robust Bellman PPO fails to recognize the invalidity of the novel output and also demonstrates negligible differences in token-level advantage estimations. Additionally, it overestimates the advantage of the final output component. This observation partially reflects the limited generalization capability of these two methods when handling novel response generations.